rowing with the current, a canoeist paddled 21 mi in 3 hr. Against the current, the canoeist could paddle only 9 mi in the same amount of time. find the rate of the canoeist in calm water and the rate of the current.

Answer by mananth(16946)

canoe speed =x mph
current speed =y mph
against current 3 hours
with current 3 hours

Distance against 9 miles distance with 21 miles
t=d/r against current -
9 / ( x - y )= 3
3 ( x - y ) = 9
3 x - 3 y = 9 ....................1

21 / ( x + y )= 3
3 ( x + y ) = 21
3 x + 3 y = 21 ...............2
Multiply (1) by 1
Multiply (2) by 1
we get y
3 x + -3 y = 9
3 x + 3 = 21
6 x = 30
/ 6
x = 5 mph

plug value of x in (1) y
3 x -3 y = 9
15 -3 -15 = 9
-3 y = 9
-3 y = -6 mph
y = 2
canoe speed 5 mph
current speed 2 mph

m.ananth@hotmail.ca